OpenStudy (anonymous):

You are buying decorations for a birthday party. Six rolls of crepe paper and 20 balloons cost $11.90. After you start decorating, you need more supplies. You buy 2 more rolls of crepe paper and 15 more balloons for $6.05. How much did each roll of crepe paper cost? A) $1.15 B) $ 1.58 C) 25¢ D) 22.6¢

hero (hero):

Hint: 6c + 20b = 11.90 2c + 15b = 6.05

OpenStudy (anonymous):

What am I supposed to do with the two equations?

hero (hero):

Apply systems of equations

OpenStudy (anonymous):

What is that?

hero (hero):

6c + 20b = 11.90 2c + 15b = 6.05 c = (11.90 - 20b)/6 c = (6.05 - 15b)/2 c = c 2(11.90 - 20b = 6(6.05 - 15b) 23.80 - 40b = 36.30 - 90b 90b - 40b = 36.30 - 23.80 50b = 12.50 b = 12.50/50 b = .25 c = 1.15 So the crept paper cost $1.15

OpenStudy (anonymous):

Thank you is what you just did a system of equations?